5x-2y=-5
y-5x=3
solve by substitution

Answer by checkley71(8403)

5X-2Y=-5
-5X+Y=3 ADDING THE TWO EQUATIONS WE GET
---------
-Y=-2 OR Y=2 SUBSTITUTING 2 FOR Y IN EQUATION 1 WE GET
5X-2*2-5 OR 5X-4=-5 OR 5X=-1 OR X=-1/5 PROOF
-5(-1/5)+2=3 OR 1+2=3 OR 3=3